OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – 2x2 – 24x ?

OpenStudy (anonymous):

x³-2x²-24x=0 Common factor x x(x²-2x-24)=0 Nule factor theorem. x=0 x²-2x-24=0 Factorize (x-6)(x+4)=0 The zeros x=0 x=6 x=-4